# Complexity results and active-set identification of a derivative-free method for bound-constrained problems

### v2

@@ 3.3 Worst-case complexity
From Theorem 2, we have that $\Delta_{m(k)+1}\geq\chi(x_{m(k)})/(\sqrt{n}c_{3})$ for all $k\in\mathcal{S}_{j_{\epsilon}}$ and $\Delta_{k+1}\geq\chi(x_{k})/(\sqrt{n}c_{2})$ for all $k\in\mathcal{U}_{j_{\epsilon}}$. Since $\chi(x_{k})\geq\epsilon$ for all $k\in\{0,\dots,j_{\epsilon}-1\}$, with $\mathcal{S}_{j_{\epsilon}}\cup\mathcal{U}_{j_{\epsilon}}=\{0,\dots,j_{\epsilon}-1\}$, we get
@@
$Nf_{j_{\epsilon}}\leq 2n\left\lfloor\frac{nc_{3}^{2}(\Phi_{0}-f_{\min})}{c_{1}}\epsilon^{-2}\right\rfloor+\left\lfloor\frac{nc_{3}^{2}(f_{0}-f_{\min})}{\gamma c^{2}}\frac{\delta^{2}}{(1-\delta)^{2}}\epsilon^{-2}\right\rfloor,$
@@
where the last inequality follows from Theorem 6. Now, let us consider any iteration $k<j_{\epsilon}$ and any index $i\in\{1,\ldots,n\}$ such that the line search succeeds to produce a decrease in the objective function. For each $\alpha$ used in the extrapolation phase of the line search, we have that
@@
Note that, by definition, $m(k)$ is the last unsuccessful iteration before iteration $k$, i.e., all the iterations from $m(k)+1$ to $k$ are successful iterations. Lemma 1 guarantees that $\Delta_{k}\geq\Delta_{m(k)+1}$ for all $k\in\mathcal{S}_{j_{\epsilon}}$. Hence, from (31), it follows that
$f(y_{k}^{i}+\alpha d_{k}^{i})-f(y_{k}^{i}+(\alpha/\delta)d_{k}^{i})\geq\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}\alpha^{2}\geq\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}c^{2}\Delta_{m(k)+1}^{2}.$
@@
$f(y_{k}^{i}+\alpha d_{k}^{i})-f(y_{k}^{i}+(\alpha/\delta)d_{k}^{i})\geq\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}c^{2}\frac{\epsilon^{2}}{nc_{3}^{2}}.$
@@
$f_{0}-f_{\min}\geq Nf_{j_{\epsilon}}^{\mathcal{S}}\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}c^{2}\frac{\epsilon^{2}}{nc_{3}^{2}},$
@@
$Nf_{j_{\epsilon}}^{\mathcal{S}}\leq\left\lfloor\frac{nc_{3}^{2}(f_{0}-f_{\min})}{\gamma c^{2}}\frac{\delta^{2}}{(1-\delta)^{2}}\epsilon^{-2}\right\rfloor.$
@@ 4 Finite active-set identification
A point $x^{*}\in[l,u]$ is a stationary point of problem (1) (i.e., $\chi(x^{*})=0$) if and only if, for all $i\in\{1,\ldots,n\}$, we have that
@@
We see that $\zeta(x^{*})$ is a measure of the minimum amount of strict complementarity among the variable in $\mathscr{Z}^{+}(x^{*})$. This quantity will be used to define a neighborhood of $x^{*}$ where the active components are correctly identified, following a similar approach as in [13, 32].
Before diving into the main theorem of this section, we need some preliminary results following from the Lipschitz continuity of $\nabla f$. Recalling Defintion 1, using standard arguments (see, e.g., [31]) one can prove that, for all $x\in[l,u]$, we have
@@
That is the line search in Algorithm 2 fails when using the direction $\operatorname{sign}(\nabla_{i}f(y_{k}^{i}))e_{i}$ with any stepsize $0<\alpha\leq\nu^{i}_{k}$. So, if the line search returns $\alpha_{k}^{i}>0$, necessarily $d_{k}^{i}=-\operatorname{sign}(\nabla_{i}f(y_{k}^{i}))e_{i}$. ∎
Now, we are ready to state the main result of this section, establishing finite active-set idenfitication of Algorithm 1.
@@
where $\alpha_{\text{max}}$ is the largest feasible stepsize along the direction $d^{i}_{k}$ at $y^{i}_{k}$. So, if $\alpha_{\text{max}}=0$, then $\alpha_{k}^{i}=0$ and, using (44), we have $(y_{k}^{i+1})_{i}=(y_{k}^{i})_{i}=l_{i}$, thus proving (37). If $\alpha_{\text{max}}>0$, using Proposition 5, (44) and (45), it follows that a sufficient decrease of $f$ along $d^{i}_{k}$ is obtained with the first stepsize $\bar{\alpha}$ used in the line search, that is, the condition at line 4 of Algorithm 2 is satisfied. Now, consider the extrapolation phase in the line search procedure, that is, lines 12–15 of Algorithm 2. Recalling (44), each stepsize $\eta=\min\{\alpha/\delta,\alpha_{\text{max}}\}$ is such that $\alpha\leq\eta\leq(y^{i}_{k})_{i}-l_{i}$, that is,
$0\leq\eta-\alpha\leq(y^{i}_{k}+\alpha d^{i}_{k})_{i}-l_{i}.$
@@
$0\leq\eta-\alpha\leq(y^{i}_{k}+\alpha d^{i}_{k})_{i}-l_{i}\leq\frac{2\nabla_{i}f(y^{i}_{k}+\alpha d^{i}_{k})}{L_{i}+2\gamma}$
for every stepsize $\eta$ used in the extrapolation. Then, using Proposition 5 with $x=y^{i}_{k}+\alpha d^{i}_{k}$, $s=\eta-\alpha$ and $d^{i}_{k}$ as in (44), it follows that
$f(y^{i}_{k}+\eta d^{i}_{k})\leq f(y^{i}_{k}+\alpha d^{i}_{k})-\gamma(\eta-\alpha)^{2}.$

### v3

@@ 3.3 Worst-case complexity
From Theorem 2, we have that $\Delta_{m(k)+1}\geq\chi(x_{m(k)})/(\sqrt{n}c_{3})$ for all $k\in\mathcal{S}_{j_{\epsilon}}$ and $\Delta_{k+1}\geq\chi(x_{k})/(\sqrt{n}c_{3})$ for all $k\in\mathcal{U}_{j_{\epsilon}}$. Since $\chi(x_{k})\geq\epsilon$ for all $k\in\{0,\dots,j_{\epsilon}-1\}$, with $\mathcal{S}_{j_{\epsilon}}\cup\mathcal{U}_{j_{\epsilon}}=\{0,\dots,j_{\epsilon}-1\}$, we get
@@
$Nf_{j_{\epsilon}}\leq 2n\left\lfloor\frac{nc_{3}^{2}(\Phi_{0}-f_{\min})}{c_{1}}\epsilon^{-2}\right\rfloor+\left\lfloor\frac{nc_{3}^{2}(f_{0}-f_{\min})}{\gamma c^{2}}\max\Bigg{\{}1,\Biggl{(}\frac{\delta}{1-\delta}\Biggr{)}^{2}\Bigg{\}}\epsilon^{-2}\right\rfloor,$
@@
where the last inequality follows from Theorem 6. Now, let us consider any iteration $k<j_{\epsilon}$ and any index $i\in\{1,\ldots,n\}$ such that the line search succeeds to produce a decrease in the objective function. For each $\alpha$ used in the extrapolation phase of the line search, we have that either
$f(y_{k}^{i})-f(y_{k}^{i}+\alpha d_{k}^{i})\geq\gamma\alpha^{2}\geq\gamma c^{2}\Delta_{k}^{2},$ (35)
or
@@
Note that, by definition, $m(k)$ is the last unsuccessful iteration before iteration $k$, i.e., all the iterations from $m(k)+1$ to $k$ are successful iterations. Lemma 1 guarantees that $\Delta_{k}\geq\Delta_{m(k)+1}$ for all $k\in\mathcal{S}_{j_{\epsilon}}$. Hence, from (35) and (36), it follows that
$f(y_{k}^{i})-f(y_{k}^{i}+\alpha d_{k}^{i})\geq\gamma\alpha^{2}\geq\gamma c^{2}\Delta_{m(k)+1}^{2}\geq\gamma\min\Bigg{\{}1,\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}\Bigg{\}}c^{2}\Delta_{m(k)+1}^{2}$
$f(y_{k}^{i}+\alpha d_{k}^{i})-f(y_{k}^{i}+(\alpha/\delta)d_{k}^{i})\geq\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}\alpha^{2}\geq\gamma{\min\Bigg{\{}1,\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}\Bigg{\}}}c^{2}\Delta_{m(k)+1}^{2}.$
@@
$f(y_{k}^{i}+\alpha d_{k}^{i})-f(y_{k}^{i}+(\alpha/\delta)d_{k}^{i})\geq\gamma\min\Bigg{\{}1,\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}\Bigg{\}}c^{2}\frac{\epsilon^{2}}{nc_{3}^{2}}.$
@@
$f_{0}-f_{\min}\geq Nf_{j_{\epsilon}}^{\mathcal{S}}\gamma{\min\Bigg{\{}1,\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}\Bigg{\}}}c^{2}\frac{\epsilon^{2}}{nc_{3}^{2}},$
@@
$Nf_{j_{\epsilon}}^{\mathcal{S}}\leq\left\lfloor\frac{nc_{3}^{2}(f_{0}-f_{\min})}{\gamma c^{2}}{\max\Bigg{\{}1,\Biggl{(}\frac{\delta}{1-\delta}\Biggr{)}^{2}\Bigg{\}}}\epsilon^{-2}\right\rfloor.$
@@ 4 Finite active-set identification
A point $x^{*}\in[l,u]$ is said to be a stationary point of problem (1) (i.e., $\chi(x^{*})=0$) if, for all $i\in\{1,\ldots,n\}$, we have that
@@
We see that $\zeta(x^{*})$ is a measure of the minimum amount of strict complementarity among the variables in $\mathscr{Z}^{+}(x^{*})$. This quantity will be used to define a neighborhood of $x^{*}$ where the active components are correctly identified, following a similar approach as in [13, 32].
Before diving into the main theorem of this section, we need some preliminary results following from the Lipschitz continuity of $\nabla f$. Recalling Definition 1, using standard arguments (see, e.g., [31]) one can prove that, for all $x\in[l,u]$, we have
@@
Thus, the line search in Algorithm 2 fails when using the direction $\operatorname{sign}(\nabla_{i}f(y_{k}^{i}))e_{i}$ with any stepsize $0<\alpha\leq\nu^{i}_{k}$. So, if the line search returns $\alpha_{k}^{i}>0$, necessarily $d_{k}^{i}=-\operatorname{sign}(\nabla_{i}f(y_{k}^{i}))e_{i}$. ∎
Now, we are ready to state the main result of this section, establishing finite active-set identification of Algorithm 1.
@@
where $\alpha_{\text{max}}$ is the largest feasible stepsize along the direction $d^{i}_{k}$ at $y^{i}_{k}$. So, if $\alpha_{\text{max}}=0$, then $\alpha_{k}^{i}=0$ and, using (49), we have $(y_{k}^{i+1})_{i}=(y_{k}^{i})_{i}=l_{i}$, thus proving (42). If $\alpha_{\text{max}}>0$, using Proposition 5, (49) and (50), it follows that a sufficient decrease of $f$ along $d^{i}_{k}$ is obtained with the first stepsize $\bar{\alpha}$ used in the line search, that is, the condition at line 4 of Algorithm 2 is satisfied. Now, consider the extrapolation phase in the line search procedure, that is, lines 12–15 of Algorithm 2. Recalling (49), each stepsize $\omega=\min\{\alpha/\delta,\alpha_{\text{max}}\}$ is such that $\alpha\leq\omega\leq(y^{i}_{k})_{i}-l_{i}$, that is,
$0\leq{\omega}-\alpha\leq(y^{i}_{k}+\alpha d^{i}_{k})_{i}-l_{i}.$
@@
$0\leq{\omega}-\alpha\leq(y^{i}_{k}+\alpha d^{i}_{k})_{i}-l_{i}\leq\frac{2\nabla_{i}f(y^{i}_{k}+\alpha d^{i}_{k})}{L_{i}+2\gamma}$
for every stepsize $\omega$ used in the extrapolation. Then, using Proposition 5 with $x=y^{i}_{k}+\alpha d^{i}_{k}$, $s=\omega-\alpha$ and $d^{i}_{k}$ as in (49), it follows that
$f(y^{i}_{k}+{\omega}d^{i}_{k})\leq f(y^{i}_{k}+\alpha d^{i}_{k})-\gamma({\omega}-\alpha)^{2}.$